the difference between 1/2 of a number and 1/3 of a number is 9. Find the number. How do I write this?
Answer by mananth(16949)

let the number be x
half the number is x/2
..
1/3 the number = x/3
..
x/2 - x/3 = 9
LCD = 6
(3x-2x)/6 = 9
x/6 =9
* my 6
x=54
...
m.ananth@hotmail.ca